# WebCiteS: Attributed Query-Focused Summarization on Chinese Web Search Results with Citations

### v1

@@ Abstract
Enhancing the attribution in large language models (LLMs) is a crucial task. One feasible approach is to enable LLMs to cite external sources that support their generations. However, existing datasets and evaluation methods in this domain still exhibit notable limitations. In this work, we formulate the task of attributed query-focused summarization (AQFS) and present WebCiteS, a Chinese dataset featuring 7k human-annotated summaries with citations. WebCiteS derives from real-world user queries and web search results, offering a valuable resource for model training and evaluation. Prior works in attribution evaluation do not differentiate between groundedness errors and citation errors. They also fall short in automatically verifying sentences that draw partial support from multiple sources. We tackle these issues by developing detailed metrics and enabling the automatic evaluator to decompose the sentences into sub-claims for fine-grained verification. Our comprehensive evaluation of both open-source and proprietary models on WebCiteS highlights the challenge LLMs face in correctly citing sources, underscoring the necessity for further improvement.
@@
Haolin Deng1111  Chang Wang2222  Xin Li2222  Dezhang Yuan2222  Junlang Zhan2222 Tianhua Zhou2222Jin Ma3333Jun Gao1111Ruifeng Xu1111thanks: Corresponding author $1$Harbin Institute of Technology, Shenzhen, China $2$Tencent Inc. $3$University of Science and Technology of China hldeng028@gmail.com, chasonwang@tencent.com, xuruifeng@hit.edu.cn
@@ results
Firstly, most existing datasets are deficient in high-quality citation annotations. For instance, the ALCE benchmark (Gao et al., 2023b) compiles three question-answering datasets (Fan et al., 2019; Stelmakh et al., 2022; Amouyal et al., 2023) without providing citations in the reference answers, limiting its utility for model training. In contrast, WebGLM  (Liu et al., 2023b) prompts InstructGPT (Ouyang et al., 2022) to generate training data with citations. It controls the citation quality via a sample filtering method which calculates the Rouge Lin (2004) score between the answers and their citations. However, this method focuses on lexical similarity rather than logical entailment, and thus could not precisely measure attribution.
Secondly, current evaluation methods are insufficient to thoroughly assess attribution. Prior works only inspect if the generations are supported by their citations Liu et al. (2023a); Gao et al. (2023b); Liu et al. (2023b) without checking all the documents provided in the input context. However, instances of unsupported generations may result from both failing to correctly cite supporting documents and failing to be grounded in all input documents. Differentiating these two types of errors is crucial for system optimization. Moreover, existing automatic evaluation (Gao et al., 2023b) solely relies on off-the-shelf natural language inference (NLI) models which only recognize entailment (full support) and overlook sentences with multiple sub-claims drawing partial support from different sources. Such complexities are common in real-world scenarios (Chen et al., 2023; Kamoi et al., 2023) and are indicative of a strong capability of synthesizing information across various sources.
To address the above limitations, we present WebCiteS, a Chinese dataset for Attributed Query-Focused Summarization (AQFS). As shown in Figure 1, given a query and the retrieved documents, AQFS aims to summarize all pertinent information from the documents with in-line citations to make the summary attributable. Our dataset is built upon real-world user queries and search results from Sogou, a widely used Chinese web search engine.111www.sogou.com We employ human efforts to ensure the quality of summaries and citations. Table 1 compares WebCiteS and relevant datasets.
@@
[…]
The dataset and code will be made available at https://github.com/HarlynDN/WebCiteS.
@@ 2.2 Data Collection
We collected 40,000 unique, anonymized user queries from Sogou, a widely used Chinese web search engine. After initial refinement, we retained 18,500 non-trivial queries and retrieved five web pages for each query. The snippets of web pages returned by the search engine were expanded to a maximum of 250 characters to serve as the documents for the annotation process.
@@ 2.3 Human-LLM Collaborative Annotation
Lastly, human annotators chose the preferred summary among the two LLM-generated candidates, refined its quality, and annotated citations. The chosen summary underwent thorough inspection with non-essential and redundant parts removed. Annotators would cite all supporting documents for verification-worthy sentences, correct groundedness and coherence errors, and supplement missing information. Offering multiple candidate summaries aimed to avoid limiting annotators to a single, potentially lower-quality option and to merge the strengths of different options, thereby improving the quality of the final summary. The core statistics are present in Table 2. More details of the dataset can be found in Appendix A.
@@
# Citations / ${}^{\dagger}$ per Sent. 1.55 / 1.62
Table 2: Core statistics of the WebCiteS dataset. Full content length and snippet length refer to the average length of a single search result (web page). ${}^{\dagger}$ Sub-claims of sentences are extracted by ChatGPT (Section 4.2).
@@ 3.1 Evaluating Summarization Utility
We apply $\psi$ to extract all sub-claims from the system summary and calculate the fraction of them being entailed by the reference summary using $\phi$. This metric could measure the accuracy and relevancy of system summaries.
@@ 3.2 Evaluating Attribution
Figure 3 shows the framework of attribution evaluation. Compared to Gao et al.’s (2023b) automatic methods, our method (1) considers citation mask, (2) incorporates the claim-split model for partial support detection, and (3) distinguishes groundedness and citation quality. See in Appendix B.2 for more discussions on attribution metrics.
@@ 4.1 Performance of the NLI model
We evaluate the performance of different open-source NLI models in predicting human-annotated citations in WebCiteS. We finally choose an mT5 model (Xue et al., 2021) fine-tuned on multilingual NLI tasks as $\phi$ for our evaluator,444huggingface.co/alan-turing-institute/mt5-large-finetuned-mnli-xtreme-xnli since it achieves the highest accuracy of 82.3% among all models. See Appendix C for more details in NLI models.
@@ 4.2 Performance of the Claim-Split Model
Redundancy, which measures if the model generates redundant sub-claims of the source sentence. Two sub-claims $c_{i}$ and $c_{j}$ in $\psi(s)$ are deemed redundant if they entail each other: $\phi(c_{i},c_{j})=1$ and $\phi(c_{j},c_{i})=1$. Based on this, we could eliminate the redundancy of $\psi(s)$: if multiple sub-claims are redundant, we only keep the first one and remove the rest. The resulting set is denoted as $\psi^{*}(s)$. Finally, the metric is computed as:
@@
# Splits, which is the average count of non-redundant sub-claims per sentence. This could reflect the granularity of model outputs, as a lower value indicates that the model may not effectively separate some sub-claims in the source sentence.
Correctness, which is the fraction of sub-claims being entailed by the source sentence using the NLI model $\phi$, as we assume that all correct sub-claims should be entailed by the source sentence.
Completeness, which is a binary value measuring if the source sentence is entailed by the concatenation of all sub-claims. If not, some essential sub-claims may be missing in the model outputs.
Table 3 shows the evaluation result. We first notice that ChatGPT and the fine-tuned mT5 models are consistent in # splits which reflects the decomposition granularity. Moreover, the mT5 models slightly outperform ChatGPT in correctness and completeness. The primary reason is that ChatGPT occasionally unfollow the task instruction (Zhou et al., 2023). Therefore, we select the fine-tuned mT5-large model as $\psi$ for our evaluator. See Appendix D for more details in claim-split models.
@@ 5.1 Main Results
[…]
@@ 5.2 Results in the Long-Context Setting
[…]
@@
[…]
@@ 6 Related Work
Relevant datasets.
Most existing datasets in query-focused multi-document summarization and long-form question answering do not consider attribution in the task setup (Fan et al., 2019; Fabbri et al., 2019; Boni et al., 2021; Stelmakh et al., 2022; Bolotova et al., 2023), while recent initiatives (Qin et al., 2023; Liu et al., 2023b; Gao et al., 2023b) exhibit limitations in citation annotation and evaluation methods detailed in Section 1 and Table 1.
@@ 7 Conclusion
In this work, we formulate the task of attributed query-focused summarization (AQFS) and present WebCiteS, a high-quality dataset derived from real-world user queries and search results. We propose a comprehensive evaluation framework on summarization utility and attribution. We also optimize the automatic evaluator to make it better aligned with human citations. Our evaluations on both open-source and proprietary LLMs underscore the unsolved challenge of attribution, especially in the long-context setting with fine-grained documents, highlighting the need for further improvement. We will open-source the dataset and code to facilitate future explorations in this crucial field.
@@ A.1 More Details of Data Annotation
After 40,000 raw queries were gathered, we adopted a rule-based system to remove common trivial queries such as looking for synonyms, spelling, or translation of a word. We also filter out queries seeking health and medicine advice, since these scenarios are of high risks and hard to judge without professional expertise. Through this process, we obtain 18,500 filtered queries. To ensure the quality of the data, the remaining samples were further filtered by human annotators. Specifically, a sample would be discarded if it matched the following scenarios:
@@
    If the query and documents contained inappropriate contents, such as discrimination against specific groups, encouragement of illegal activities, and controversial sentences regarding politics or religions.
@@
    If the query could not be answered by the retrieved documents.
@@ Appendix C Experiments on NLI Model
[…]

### v2

@@ Abstract
Enhancing the attribution in large language models (LLMs) is a crucial task. One feasible approach is to enable LLMs to cite external sources that support their generations. However, existing datasets and evaluation methods in this domain still exhibit notable limitations. In this work, we formulate the task of attributed query-focused summarization (AQFS) and present WebCiteS, a Chinese dataset featuring 7k human-annotated summaries with citations. WebCiteS derives from real-world user queries and web search results, offering a valuable resource for model training and evaluation. Prior works in attribution evaluation do not differentiate between groundedness errors and citation errors. They also fall short in automatically verifying sentences that draw partial support from multiple sources. We tackle these issues by developing detailed metrics and enabling the automatic evaluator to decompose the sentences into sub-claims for fine-grained verification. Our comprehensive evaluation of both open-source and proprietary models on WebCiteS highlights the challenge LLMs face in correctly citing sources, underscoring the necessity for further improvement. [^1] The dataset and code are released under Apache License 2.0 at https://github.com/HarlynDN/WebCiteS
@@
Haolin Deng1111thanks: This work was done during the internship at Tencent. Chang Wang2222  Xin Li2222  Dezhang Yuan2222  Junlang Zhan2222 Tianhua Zhou2222Jin Ma3333Jun Gao111122footnotemark: 2Ruifeng Xu1111thanks: Corresponding authors $1$Harbin Institute of Technology, Shenzhen, China $2$Tencent $3$University of Science and Technology of China hldeng028@gmail.com, dezhangyuan@tencent.com, xuruifeng@hit.edu.cn
@@ results
Firstly, most existing datasets are deficient in high-quality citation annotations.
For instance, the ALCE benchmark (Gao et al., 2023b) compiles three question-answering datasets (Fan et al., 2019; Stelmakh et al., 2022; Amouyal et al., 2023) without providing citations in the reference answers, limiting its utility for model training. In contrast, WebGLM  (Liu et al., 2023b) prompts InstructGPT (Ouyang et al., 2022) to generate training data with citations. It controls the citation quality via a sample filtering method which calculates the ROUGE Lin (2004) score between the answers and their citations. However, this method focuses on lexical similarity rather than logical entailment, and thus could not precisely measure attribution.
Secondly, current evaluation methods are insufficient to thoroughly assess attribution.
Prior works only inspect if the generations are supported by their citations Liu et al. (2023a); Gao et al. (2023b); Liu et al. (2023b) without checking all the documents provided in the input context. However, instances of unsupported generations may result from both failing to correctly cite supporting documents and failing to be grounded in all input documents. Differentiating these two types of errors is crucial for system optimization. Moreover, existing automatic evaluation (Gao et al., 2023b) solely relies on off-the-shelf natural language inference (NLI) models which only recognize entailment (full support) and overlook sentences with multiple sub-claims drawing partial support from different sources. Such complexities are common in real-world scenarios (Chen et al., 2023; Kamoi et al., 2023) and are indicative of a strong capability of synthesizing information across various sources.
To address the above limitations, we present WebCiteS, a Chinese dataset for Attributed Query-Focused Summarization (AQFS). As shown in Figure 1, given a query and the retrieved documents, AQFS aims to summarize all pertinent information from the documents with in-line citations to make the summary attributable. Our dataset is built upon real-world user queries and search results from Sogou, a widely used Chinese web search engine.222www.sogou.com We employ human efforts to ensure the quality of summaries and citations. Table 1 compares WebCiteS and relevant datasets.
@@
Our evaluation of both open-source and proprietary models on WebCiteS reveals the following key findings: (1) contextual grounding of generations does not guarantee the avoidance of citation errors, indicating the challenge of explicit attribution in all the tested LLMs; (2) although supervised fine-tuning improves both groundedness and citation quality, the top-performing model only reaches a citation F1 score of 76.1% with about 20% of sentences not being fully supported by their citations, underscoring the need for further optimization; (3) models perform worse when summarizing full content of web pages rather than shorter snippets, showing that LLMs are less effective at synthesizing and attributing information in the longer context; (4) making documents more fine-grained leads to poorer attribution results, highlighting the difficulty LLMs face in pinpointing the exact supporting evidence within the context.
@@ 2.2 Data Collection
We collected 40,000 unique, anonymized user queries from Sogou, a widely used Chinese web search engine, encapsulating a diverse range of real-world questions. After initial refinement, we retained 18,500 non-trivial queries and retrieved five web pages for each query.333 Common trivial queries include ones that look for word synonyms, text translation, celebrity birth dates, etc. During annotation, we found the top five search results were adequate to address most queries. The snippets of web pages returned by the search engine were expanded to a maximum of 250 characters to serve as the documents for the annotation process.
@@ 2.3 Human-LLM Collaborative Annotation
Lastly, human annotators chose the preferred summary among the two LLM-generated candidates, refined its quality, and annotated citations. The chosen summary underwent thorough inspection with non-essential and redundant parts removed. Annotators would cite all supporting documents for verification-worthy sentences, correct groundedness and coherence errors, and supplement missing information. Offering multiple candidate summaries aimed to avoid limiting annotators to a single, potentially lower-quality option and to merge the strengths of different options, thereby improving the quality of the final summary.
@@
# Citations / Sub-claims per Sent. 1.55 / 1.62
Table 2: Core statistics of the WebCiteS dataset. Full content length and snippet length refer to the average length of a single search result (web page). Sub-claims of sentences are extracted by ChatGPT (Section 4.2).
Quality control.
The core statistics of WebCiteS are present in Table 2. Moreover, we conduct the following analysis on the quality of the dataset:
Are the retrieved documents useful to the queries?
Though we did not ask annotators to explicitly label the relevance score for each document, the human-extracted information from stage 1 could reflect how useful the document is to the query. After annotation, we find that 87.2% of the documents have human extraction, while the average length of extracted segments per document is 93.4 characters. This indicates that the majority of retrieved documents are helpful to the query.
How much manual refinement is made on candidate summaries?
The average Levenshtein distance between the human-refined summary and the human-preferred candidate summary is 74.1 (we only count summary edits, excluding citation annotation). This suggests that the quality of candidate summaries were generally judged imperfect by the human annotators.
Overlap of web pages.
See Appendix A for more details of the dataset.
@@ 3.1 Evaluating Summarization Utility
We apply $\psi$ to extract all sub-claims from the system summary and calculate the fraction of them being entailed by the reference summary using $\phi$. This metric could measure the accuracy and relevance of system summaries.
@@ 3.2 Evaluating Attribution
Figure 3 shows the framework of attribution evaluation. Compared to Gao et al.’s (2023b) automatic methods, our method (1) considers citation mask, (2) incorporates the claim-split model for partial support detection, and (3) distinguishes groundedness and citation quality. See Appendix B.2 for more discussions on attribution metrics.
@@ 4.1 Performance of the NLI model
We evaluate the performance of different open-source NLI models in predicting human-annotated citations in WebCiteS. We finally choose an mT5 model (Xue et al., 2021) fine-tuned on multilingual NLI tasks as $\phi$ for our evaluator,666huggingface.co/alan-turing-institute/mt5-large-finetuned-mnli-xtreme-xnli since it achieves the highest accuracy of 82.3% among all models. See Appendix C for more details about NLI models.
@@ 4.2 Performance of the Claim-Split Model
Redundancy.
It measures if the model generates redundant sub-claims of the source sentence. Two sub-claims $c_{i}$ and $c_{j}$ in $\psi(s)$ are deemed redundant if they entail each other: $\phi(c_{i},c_{j})=1$ and $\phi(c_{j},c_{i})=1$. Based on this, we could eliminate the redundancy of $\psi(s)$: if multiple sub-claims are redundant, we only keep the first one and remove the rest. The resulting set is denoted as $\psi^{*}(s)$. Finally, the metric is computed as:
@@
# Splits.
It is defined as the average count of non-redundant sub-claims per sentence, which could reflect the granularity of model outputs, as a lower value indicates that the model may not effectively separate some sub-claims in the source sentence.
Correctness.
It is defined as the fraction of sub-claims being entailed by the source sentence using the NLI model $\phi$, as we assume that all correct sub-claims should be entailed by the source sentence.
Completeness.
It is a binary value measuring if the source sentence is entailed by the concatenation of all sub-claims using the NLI model $\phi$. If not, some essential sub-claims may be missing in the model outputs.
Table 3 shows the evaluation result. We first notice that ChatGPT and the fine-tuned models are consistent in # splits which reflects the decomposition granularity. Moreover, the fine-tuned models slightly outperform ChatGPT in correctness and completeness. The primary reason is that ChatGPT occasionally unfollow the task instruction (Zhou et al., 2023). Therefore, we select the fine-tuned mT5-large model as $\psi$ for our evaluator. See Appendix D for more details about claim-split models.
@@ 5.1 Main Results
We first adopt the default setting where each sample consists of five snippets of web pages as documents. The experimental results are present in Table 5. In general, we observe a large variance of summary lengths across models. For FSP, ChatGPT and GPT-4 outperform all open-source LLMs on both claim F1 scores and citation F1 scores. However, even GPT-4 exhibits unignorable attribution errors: only 72% of its citations are correct and only 71% of supporting documents are cited. Moreover, only 76% of its generated sentences are fully supported by their respective citations, and only 81% of them are grounded in the input context. For SFT, we observe that smaller pre-trained models such as mT5 significantly lag behind open-source LLMs on both summarization utility and citation quality. Although the fine-tuned mT5-Large model achieves the best groundedness (reflected by the highest ACS score of 90.3%), we find it is primarily due to the model’s tendency to copy the input text rather than summarize the content pertinent to the query, which leads to the suboptimal claim F1 scores. Our additional findings include:
@@ 5.2 Results in the Long-Context Setting
We observe an overall decline in both summarization utility and attribution with full content instead of snippets. We also visualize the performance variance over the context length in Figure 4, which shows a decline in claim F1, citation F1, and AIS scores as the context length extends. This pattern indicates that extending context length challenges models’ ability to synthesize pertinent information and correctly cite sources.
@@
In practice, if the cited documents are too long, it is challenging to verify the model generations, and we expect the models could cite more specific segments of information. Therefore, we also vary the maximum document length when chunking the web pages and investigate the impact of document (citation) granularity. As shown in Table 6, reducing max document length from 512 to 256, without changing the total input evidence, drastically lowers citation F1 and AIS scores for all three models. This performance reduction reveals the difficulty LLMs face in precisely pinpointing the exact supporting evidence within the context.
@@ 6 Related Work
Relevant datasets and benchmarks.
Query-focused multi-document summarization (QF-MDS) aims to summarize multiple documents driven by specific queries (Tombros and Sanderson, 1998; Li and Li, 2013; Roy and Kundu, 2023). Similarly, long-form question answering (LFQA) focuses on producing detailed answers, often utilizing external sources (Krishna et al., 2021). Most existing datasets in both tasks do not consider attribution in the setup (Fan et al., 2019; Fabbri et al., 2019; Boni et al., 2021; Stelmakh et al., 2022; Bolotova et al., 2023). Bohnet et al. (2022) propose the attributed question answering (AQA) benchmark where the system must output the answer alongside a piece of evidence. However, long-form responses usually require citing multiple sources of evidence. Recent initiatives (Qin et al., 2023; Liu et al., 2023b; Gao et al., 2023b) in this dimension exhibit limitations in citation annotation and evaluation methods detailed in Section 1 and Table 1.
@@ 7 Conclusion
In this work, we formulate the task of attributed query-focused summarization (AQFS) and present WebCiteS, a high-quality dataset derived from real-world user queries and search results. We propose a comprehensive evaluation framework on summarization utility and attribution. Notably, we highlight two fine-grained dimensions of attribution: groundedness and citation quality. We further enhance the framework with a carefully-designed automatic evaluator, and validate its substantial agreement with human annotators. Finally, we evaluate both open-source and proprietary LLMs extensively on WebCiteS to underscore the unsolved challenge of attribution, especially in the long-context setting with fine-grained documents. We believe WebCiteS could facilitate more future explorations in attributed language models.
@@
Ethics Statement
Since WebCiteS is built upon real user queries, we have taken strict measures to address privacy issues. We only sampled anonymized queries from the search log without collecting any other information such as user identifiers. All queries were shuffled and not present in time order, making it impossible to obtain individual search history from the dataset. Lastly, during annotation, we asked annotators and quality inspectors to pay attention to discard any query with potential privacy issues.
Moreover, we have endeavored to eliminate all inappropriate content from our corpus. Firstly, we adopt an internal commercial tool to automatically detect and discard queries with improper intentions. Secondly, the commercial search engine used in this work has taken content quality and safety into account during web page retrieval and ranking, so it is unlikely that the top five search results would contain dubious or harmful material. Thirdly, human annotators were asked to discard any samples with inappropriate content (see Appendix A). Such a manual inspection process, despite being essential to enhance the dataset quality, would inevitably expose potential risks to the annotators themselves. We attempted to minimize those risks by automatically filtering most inappropriate content with the commercial tool and search engine before annotation. In practice, of all samples discarded by annotators, only less than 5% fell into the category of inappropriate content. Meanwhile, we alerted annotators of the potential risks in advance by including a cautionary note in the annotation instruction and allowed them to skip any sample that made them uncomfortable.
Finally, this work focuses on attribution rather than factuality: even if a response is fully supported by the evidence it attributes, it is not guaranteed to be factual since the evidence itself might contain factual errors or become outdated. Just as previous works (Bohnet et al., 2022; Rashkin et al., 2023) point out, the judge of factuality, especially in open domains, is extremely difficult. During the construction of WebCiteS, though we requested annotators to discard samples with highly questionable materials, we did not assume that their professional expertise could cover all fields in the corpus. Therefore, we emphasize that future works should be cautious about treating the annotated summaries in WebCiteS as "facts".
@@ A.1 More Details of Data Annotation
After 40,000 raw queries were gathered, we adopted a rule-based system to remove common trivial queries. We also filter out queries seeking health and medicine advice, since these scenarios are of high risks and hard to judge without professional expertise. Through this process, we obtain 18,500 filtered queries. To ensure the quality of the data, the remaining samples were further filtered by human annotators. Specifically, a sample would be discarded if it matched the following scenarios:
@@
    If the query and documents contained inappropriate content, such as personal information, prejudice or bias against specific groups, and controversial topics.
@@
    If the query could not be answered by the retrieved documents, or the reliability of certain documents was highly questionable to the annotators.
@@ Appendix C Experiments on NLI Model
Source Type Max Doc Length # Docs Len. Self- Bleu ↓ Prec. Rec. F1 Prec. Rec. F1 AIS ACS